# Hallucination Detection in Foundation Models for Decision-Making: A Flexible Definition and Review of the State of the Art

### v1

@@ 5.1 Metrics
5.1.1 Language Similarity
BERTScore Zhang et al. (2020)
Given a pair of responses, BERTScore computes the BERT Devlin et al. (2019) embeddings of the sentences and calculates their cosine similarity.
BARTScore Yuan et al. (2021)
Using a pre-trained BART model, which provides access to generated token probabilities, BARTScore sums over the log probability of each token generated while conditioning on context and previously output tokens. Essentially, BARTScore attempts to predict the quality of a generated text using BART as a proxy model.
SummaC Laban et al. (2022)
@@ 5.1.1 Language Similarity
Semantic Uncertainty Kuhn et al. (2023)
One common method of measuring uncertainty of a model’s many generations is computing its entropy over all generated token probabilities. However, in cases where multiple sentences have the same semantic meaning but output different entropies, the aggregated measurement is not representative of the true uncertainty of the model. Kuhn et al. (2023) tackle this problem by clustering sentences into semantic classes and summing entropies of sentences from the same class together.
5.1.2 Object Detection
CHAIR Rohrbach et al. (2018)
@@ 5.1.2 Object Detection
POPE Li et al. (2023c)
The authors recognize that different instructions prompting for a description of an image may lead to different responses from the model with the same semantic meaning. In this case, CHAIR gives different scores to both descriptions although they are alike. Instead, their proposed metric, POPE, asks binary questions about the existence of in-domain and out-of-domain objects in the image, which leads to more a more stable metric across different outputs.
5.2 Offline Datasets
@@ 5.2.1 Driving
DriveGPT4 Xu et al. (2024)
The authors augment BDD-X into a QA dataset consisting of questions that ask about the current action of the vehicle, reasoning behind the action, and predicting future control signals. To incorporate other questions a user might ask about the vehicle, surroundings, and other miscellaneous queries, they prompt ChatGPT to generate further questions. In total, the DriveGPT4 dataset contains $56$K samples.
nuScenes Caesar et al. (2020)
@@
Refer-KITTI Wu et al. (2023a)
While Talk2Car is a pioneering work for object referral in real driving scenes through natural language, each annotated instruction only refers to one object. As such, Wu et al. (2023a) propose a new task definition, referring multi-object tracking (RMOT), which attempts to predict all objects that are referred to within a natural language input. They augment the KITTI driving dataset Geiger et al. (2012) with labeled $2$D bounding boxes around objects that are referenced within a text prompt for $6.5$K images.
NuPrompt Wu et al. (2023b)
NuPrompt is another RMOT-based benchmark, but applied to nuScenes and with $3$D bounding box labels. It includes $35$K languages prompts, with most prompts referring to anywhere between one and ten objects in a scene.
DRAMA Malla et al. (2023)
The authors argue that, while several datasets exist for anomaly detection or identification on roads, there is a gap in explaining the reason for categorizing an object as being risky, i.e., objects the model should pay attention to, like crosswalks, pedestrians, and traffic lights. As such, DRAMA is a benchmark tackling identification of risky objects in a driving scene conditioned on natural language. Ding et al. (2023) extend DRAMA to further include suggestions on actions the ego vehicle can take to minimize risk, but the dataset is not public at this time.
NuInstruct Ding et al. (2024)
@@
DriveLM Sima et al. (2023)
The authors of DriveLM curate a similar comprehensive dataset from nuScenes and the CARLA driving simulator Dosovitskiy et al. (2017) with open-ended and factual questions about importance rankings of nearby vehicles, planning actions, detecting lanes, and more.
Driving with LLMs Chen et al. (2023c)
The authors collect a text-based QA dataset from a proprietary driving simulator, generated from ChatGPT with ground-truth observations (e.g., relative locations of detected vehicles, ego vehicle control actions, etc.) from the simulator.
5.2.2 Code Generation and Robotics
HumanEval Chen et al. (2021)
@@ 5.2.2 Code Generation and Robotics
RoboCodeGen Liang et al. (2023)
The authors build a new code generation benchmark specifically for robot tasks with $37$ functions focused on spatial reasoning, geometric reasoning, and controls.
Language-Table Lynch et al. (2023)
@@
5.2.3 Question-answering
HotPotQA Yang et al. (2018)
@@ 5.2.3 Question-answering
Natural Questions Kwiatkowski et al. (2019)
@@
QreCC Anantha et al. (2021)
Separate from the information retrieval task described in benchmarks above, Anantha et al. (2021) develop a dataset, QreCC, for conversational QA. They focus on reading comprehension, passage retrieval, and question rewriting tasks, with a total of $13.7$K dialogues paired with $81$K questions.
HA-DPO Zhao et al. (2024)
Zhao et al. (2024) present a multi-model visual QA dataset of images, hallucinated descriptions, and non-hallucinated samples from the VG dataset Krishna et al. (2017).
5.3 Simulation Platforms
@@
5.3.1 Driving
HighwayEnv Leurent (2018)
The developer presents a $2$D car simulator, with driving scenarios ranging from a passing on a multi-lane highway, merging into a highway, merging and exiting from a roundabout, parking, and more. An ego vehicle can be controlled with discrete (e.g., merge left, merge right, faster, etc.) or continuous (e.g., providing an explicit acceleration command) actions.
SUMO Lopez et al. (2018)
@@ 5.3.1 Driving
CARLA Dosovitskiy et al. (2017)
CARLA is a $3$D driving simulator built on top of Unreal Engine. Existing works benchmark their methods on CARLA for perception, planning, control, and QA tasks for its realism. There is also capability to perform co-simulation with SUMO and CARLA simultaneously Wegener et al. (2008).
5.3.2 Robotics
Ravens Zeng et al. (2021)
Ravens is a $3$D manipulator robot (UR5e) simulator built with PyBullet Coumans and Bai (2016–2021) with tasks like block insertion, towers of hanoi, aligning boxes, assembling kits, etc. Each simulated task features a manipulator robot with a suction gripper sitting on a table workspace, with three camera views.
ALFWorld Shridhar et al. (2021)
Building on top of the TextWorld simulator, discussed in Section 5.3.3, ALFWorld aligns perception from the $3$D robot simulation benchmark, ALFRED Shridhar et al. (2020), with text-based, discrete actions like “MoveAhead,” “RotateLeft,” and “Open.”
ProgPrompt Singh et al. (2023)
@@ 5.3.2 Robotics
RoboEval Hu et al. (2024)
RoboEval is a general platform for checking the correctness of code generated for a robot task. It relies on a simulator, evaluator, and a set of defined tasks to perform evaluations on a simulated robot. While ProgPrompt captures more realistic scenarios in its high-fidelity $3$D simulator, RoboEval is more tuned towards verifying code efficiently.
KnowNo TableSim Ren et al. (2023)
@@
5.3.3 Other Simulators
TextWorld Côté et al. (2019)
@@ 5.3.3 Other Simulators
BabyAI Chevalier-Boisvert et al. (2019)
@@
Chevalier-Boisvert et al. (2019) present a $2$D top-down, grid-based simulator of instruction-following tasks with varying difficulty. Some tasks include simple navigation to a single goal, picking and placing objects with ambiguous references, and instructions that implicitly require multi-step reasoning to complete. The simulator provides a partial observation of the space near the agent at every timestep.
MineDojo Fan et al. (2022)
@@
The developers of MineDojo attempt to create a benchmark to test the continual learning of agents in an open-world setting. They build an interface on top of Minecraft, a video game, to enable testing with diverse open-ended tasks, and provide access to an external knowledge-base of existing Minecraft tutorials and wiki discussions. MineDojo includes several thousands of tasks that are more complex that earlier works (and require multi-step reasoning). As such, task completion is judged with a learned LVLM, which acts like a human evaluator.
Smallville Park et al. (2023)
@@
The authors present a multi-agent conversational simulator where agents are controlled by language models. Users may set up agents with a defined backstory and provide instructions when desired. Each agent has access to a memory of past experiences, and generates natural language actions to go to certain areas, communicate with others, complete chores, and more.
6 Future Directions
Evaluating Methods on Decision-Making Tasks
Most hallucination detection approaches are currently evaluated in offline QA settings for information retrieval or knowledge alignment, as seen in Table 2. As foundation models are increasingly used for more complex tasks, researchers should make an effort to adapt and evaluate earlier detection/mitigation approaches that were applied to QA problems in these modern applications. Although dissimilar in practice from QA settings, planning and control problems may be formulated in such a way that enables these earlier mitigation methods to be evaluated on decision-making tasks. For example, as discussed in Section 2.1Chen et al. (2023c) treat the autonomous driving task as a QA problem, which could be naturally extended to test other QA hallucination detection methods in the same setting. This evaluation may lead to greater understanding of the general limitations of these models, as we draw parallels across diverse deployments.
Development of More Black-box Approaches
White- and grey-box detection methods may not generally be applicable in situations where the internal state or token probabilities are unavailable from the language model. Thus, we predict black-box approaches will take precedence in the near future, as state-of-the-art LVLMs like GPT-4V already prohibit access to probability outputs. However, current black-box methods are limited with simplistic sampling techniques to gauge uncertainty, and proxy models may not be representative of the true state of the model under test. Works like FLIRT (while only applied to image generation models) showcase the promise of black-box adversarial prompting approaches in generating undesirable results from models Mehrabi et al. (2023). We argue developing more aggressive black-box adversarial generative models, which explicitly optimize for producing inputs that may perturb the system outputs, is key to identifying the limits of a foundation model’s knowledge.
Pushing Models’ Generalization Capabilities
Currently, foundation models are primarily deployed to decision-making tasks that likely have some relation to its training set. For example, although complex, tasks like multi-agent communication, autonomous driving, and code generation will be present in training datasets. On the other hand, dynamic environments like robot crowd navigation require identifying nuances in pedestrian behaviors which the model may not have explicitly seen during training. Pushing the limits of foundation model deployments will allow researchers to find areas for growth in other applications.
Testing Multi-modal Models
With the explosion of LVLMs, which allow for explicit grounding of natural language and vision modalities, further exploration should be performed in evaluating their effectiveness in decision-making systems. Wen et al. (2023) take a step in the right direction towards testing black-box LVLMs in offline driving scenarios, but there is still work to be done in deploying these models in online settings. This direction can shed light on the long-standing debate of whether modular or end-to-end systems should be preferred in a particular deployment setting.
Summary
In this survey, we provide a glimpse into the progress of research into evaluating hallucinations of foundation models for decision-making problems. We begin by identifying existing usecases of foundation models in decision-making applications like autonomous driving and robotics, and find several works make note of undesired hallucinated generations in practice. By referencing works that encounter hallucinations across diverse domains, we provide a flexible definition for hallucinations that researchers can leverage, regardless of the deployment scenario in mind. Finally, we give a taxonomy of hallucination detection and mitigation approaches for decision-making problems, alongside a list of commonly used metrics, datasets, and simulators for evaluation. We find that existing methods range in varying assumptions of inputs and evaluation settings, and believe there is much room for growth in general, black-box hallucination detection algorithms for foundation models.

### v2

@@ D.1. Metrics
D.1.1. Language Similarity
BERTScore (Zhang et al., 2020)
Given a pair of responses, BERTScore computes the BERT (Devlin et al., 2019) embeddings of the sentences and calculates their cosine similarity.
BARTScore (Yuan et al., 2021)
Using a pre-trained BART model (Lewis et al., 2020a), which provides access to generated token probabilities, BARTScore sums over the log probability of each token generated while conditioning on context and previously output tokens. Essentially, BARTScore attempts to predict the quality of a generated text using BART as a proxy model.
SummaC (Laban et al., 2022)
@@ D.1.1. Language Similarity
Semantic Uncertainty (Kuhn et al., 2023)
One common method of measuring uncertainty of a model’s many generations is computing its entropy over all generated token probabilities. However, in cases where multiple sentences have the same semantic meaning but output different entropies, the aggregated measurement is not representative of the true uncertainty of the model. Kuhn et al. tackle this problem by clustering sentences into semantic classes and summing entropies of sentences from the same class together.
D.1.2. Image Captioning
CHAIR (Rohrbach et al., 2018)
@@ D.1.2. Image Captioning
POPE (Li et al., 2023b)
Li et al. recognize that different instructions prompting for a description of an image may lead to different responses from the model with the same semantic meaning. In this case, CHAIR gives different scores to both descriptions although they are alike. Instead, their proposed metric, POPE, asks binary questions about the existence of in-domain and out-of-domain objects in the image, which leads to more a more stable metric across different outputs.
D.2. Offline Datasets
@@ D.2.1. Driving
DriveGPT4 (Xu et al., 2024)
Xu et al. augment BDD-X into a QA dataset consisting of questions that ask about the current action of the vehicle, reasoning behind the action, and predicting future control signals. To incorporate other questions a user might ask about the vehicle, surroundings, and other miscellaneous queries, they prompt ChatGPT to generate further questions. In total, the DriveGPT4 dataset contains $56$K samples.
nuScenes (Caesar et al., 2020)
@@
Refer-KITTI (Wu et al., 2023a)
While Talk2Car is a pioneering work for object referral in real driving scenes through natural language, each annotated instruction only refers to one object. As such, Wu et al. propose a new task definition, referring multi-object tracking (RMOT), which attempts to predict all objects that are referred to within a natural language input. They augment the KITTI driving dataset (Geiger et al., 2012) with labeled $2$D bounding boxes around objects that are referenced within a text prompt for $6.5$K images.
NuPrompt (Wu et al., 2023b)
NuPrompt is another RMOT-based benchmark, but applied to nuScenes and with $3$D bounding box labels. It includes $35$K languages prompts, with most prompts referring to anywhere between one and ten objects.
DRAMA (Malla et al., 2023)
Malla et al. argue that, while several datasets exist for anomaly detection or identification on roads, there is a gap in explaining the reason for categorizing an object as being risky, i.e., objects the model should pay attention to, like crosswalks, pedestrians, and traffic lights. As such, DRAMA is a benchmark tackling identification of risky objects in a driving scene conditioned on natural language. Ding et al. (2023) extend DRAMA to further include suggestions on actions the ego vehicle can take to minimize risk, but the dataset is not public at this time.
NuInstruct (Ding et al., 2024)
@@
DriveLM (Sima et al., 2023)
The authors of DriveLM curate a similar comprehensive dataset from nuScenes and the CARLA driving simulator (Dosovitskiy et al., 2017) with open-ended and factual questions about importance rankings of nearby vehicles, planning actions, detecting lanes, and more.
Driving with LLMs (Chen et al., 2024b)
Chen et al. collect a text-based QA dataset from a proprietary driving simulator, generated from ChatGPT with ground-truth observations (e.g., relative locations of detected vehicles, ego vehicle control actions, etc.) from the simulator.
D.2.2. Code Generation and Robotics
HumanEval (Chen et al., 2021)
@@ D.2.2. Code Generation and Robotics
RoboCodeGen (Liang et al., 2023)
Liang et al. build a new code generation benchmark specifically for robot tasks with $37$ functions focused on spatial reasoning, geometric reasoning, and controls.
Language-Table (Lynch et al., 2023)
@@
D.2.3. Question-answering
HotPotQA (Yang et al., 2018)
@@ D.2.3. Question-answering
Natural Questions (Kwiatkowski et al., 2019)
@@
QreCC (Anantha et al., 2021)
Separate from the information retrieval task described in benchmarks above, Anantha et al. develop a dataset, QreCC, for conversational QA. They focus on reading comprehension, passage retrieval, and question rewriting tasks, with a total of $13.7$K dialogues paired with $81$K questions.
HA-DPO (Zhao et al., 2024b)
Zhao et al. present a multi-model visual QA dataset of images, hallucinated descriptions, and non-hallucinated samples from the VG dataset (Krishna et al., 2017).
D.2.4. Image and Video Generation
AVA (Murray et al., 2012)
Murray et al. tackle curating a dataset to evaluate the aesthetic value of varying quality images. AVA contains $250$K images with aesthetic ratings, textual semantic class labels, and pictographic style labels. The majority of images contain at least one semantic label, with $150$K containing at least two labels.
MSCOCO (Lin et al., 2014)
The MSCOCO dataset contains $328$K images with $2.5$M human-labeled segmentations of common objects. Each image is accompanied by five text captions. In contrast to earlier datasets like ImageNet (Deng et al., 2009) and SUN (Xiao et al., 2010), MSCOCO provides more instance labels per class and more non-iconic images containing multiple objects per scene.
LSUN (Yu et al., 2015)
As vision models continue to evolve and essentially solve existing benchmarks, new datasets need to be curated to provide additional challenges and evaluate generalization performance. Thus, Yu et al. propose a human-in-the-loop data labeling scheme to gather a dataset with one million images for ten scene and $20$ object classes, by iterating between manual labeling, overfitting a classifier, and automatic labeling. Note that the overall precision of labels is slightly lower than pure manual labeling approaches.
LAION-$400$(Schuhmann et al., 2021)
More recently, companies are collecting proprietary datasets to train large vision models, outperforming open-sourced models trained on smaller scale datasets. To narrow the gap in training data for generative models, Schuhmann et al. present an image dataset with $400$M images with corresponding metadata, CLIP (Radford et al., 2021) embeddings, and web crawling resources. Unfortunately, works have found that the training split of LAION-$400$M contains undesired content like racy imagery, slurs, stereotypes, and other biases, which should be filtered out (Saharia et al., 2022).
I2P (Schramowski et al., 2023)
The Inappropriate Image Prompts dataset attempts to measure the likelihood of image generation models to output undesired, inappropriate content given a prompt. In particular, the dataset contains $4.5$K prompts taken from real users of Stable Diffusion, which have been filtered to contain inappropriate details like hate speech, violent imagery, and criminal behavior.
DrawBench (Saharia et al., 2022)
The DrawBench benchmark contains $200$ prompts across $11$ categories to evaluate the quality of generated images on a spectrum of properties (e.g., colors, quantities, generated text). Saharia et al. use their benchmark to present $8$ samples from two different models to human raters, who choose which model they prefer qualitatively.
T2VHaluBench (Chu et al., 2024)
The developers of the Sora Detector curate a benchmark dataset for text to video hallucination detection. In total, the dataset contains $59$, $8$-second-long videos — $53$ of which contain consistency, static, or dynamic hallucinations generated by the Runway-Gen-2 (AI, 2023) and SORA (OpenAI, 2024) generative models.
D.3. Simulation Platforms
@@
D.3.1. Driving
HighwayEnv (Leurent, 2018)
Leurent presents a $2$D car simulator, with driving scenarios ranging from a passing on a multi-lane highway, merging into a highway, merging and exiting from a roundabout, parking, and more. An ego vehicle can be controlled with discrete (e.g., merge left, merge right, faster, etc.) or continuous (e.g., providing an explicit acceleration command) actions.
SUMO (Lopez et al., 2018)
@@ D.3.1. Driving
CARLA (Dosovitskiy et al., 2017)
CARLA is a $3$D driving simulator built on top of Unreal Engine. Existing works benchmark their methods on CARLA for perception, planning, control, and QA tasks for its realism. There is also capability to perform co-simulation with SUMO and CARLA simultaneously (Wegener et al., 2008).
D.3.2. Robotics
Ravens (Zeng et al., 2021)
Ravens is a $3$D manipulator robot (UR5e) simulator built with PyBullet (Coumans and Bai, 2021) with tasks like block insertion, towers of hanoi, aligning boxes, assembling kits, etc. Each simulated task features a manipulator robot with a suction gripper sitting on a table workspace, with three camera views.
ALFWorld (Shridhar et al., 2021)
Building on top of the TextWorld simulator, discussed in Appendix D.3.3, ALFWorld aligns perception from the $3$D robot simulation benchmark, ALFRED (Shridhar et al., 2020), with text-based, discrete actions like “MoveAhead,” “RotateLeft,” and “Open.”
ProgPrompt (Singh et al., 2023)
@@ D.3.2. Robotics
RoboEval (Hu et al., 2024)
RoboEval is a general platform for checking the correctness of code generated for a robot task. It relies on a simulator, evaluator, and a set of defined tasks to perform evaluations on a simulated robot. While ProgPrompt captures more realistic scenarios in its high-fidelity $3$D simulator, RoboEval is tuned towards verifying code efficiently.
KnowNo TableSim (Ren et al., 2023)
@@
D.3.3. Other Simulators
TextWorld (Côté et al., 2019)
@@ D.3.3. Other Simulators
BabyAI (Chevalier-Boisvert et al., 2019)
@@
Chevalier-Boisvert et al. present a $2$D top-down, grid-based simulator of instruction-following tasks with varying difficulty. Some tasks include simple navigation to a single goal, picking and placing objects with ambiguous references, and instructions that implicitly require multi-step reasoning to complete. The simulator provides a partial observation of the space near the agent at every timestep.
MineDojo (Fan et al., 2022)
@@
The developers of MineDojo attempt to create a benchmark to test the continual learning of agents in an open-world setting. They build an interface on top of Minecraft, a video game, to enable testing with diverse open-ended tasks, and provide access to an external knowledge-base of existing Minecraft tutorials and wiki discussions. MineDojo includes several thousands of tasks that are more complex than earlier works (and require multi-step reasoning). As such, task completion is judged with a learned LVLM, which acts like a human evaluator.
Smallville (Park et al., 2023)
@@
Park et al. present a multi-agent conversational simulator where agents are controlled by language models. Users may set up agents with a defined backstory and provide instructions when desired. Each agent has access to a memory of past experiences, and generates natural language actions to go to certain areas, communicate with others, complete chores, and more.
Appendix E Guidelines on Current Methodologies
E.1. Beware of Erroneous Hallucination Predictions
We additionally point out risks of relying on hallucination intervention algorithms and metrics that use deep learning methods at inference time. Generally speaking, deep neural networks are black boxes with few statistical guarantees and lack interpretability (Aggarwal, 2023; Dobson, 2023; Benitez et al., 1997; Taheri et al., 2021; Poggio et al., 2020; Buskulic et al., 2024). As such, these models may incorrectly detect hallucinations during deployment — effectively hallucinating themselves. In particular, engineers should take precautions when utilizing detection algorithms under method types like hidden states, attention weights, concept probabilities, analyzing samples, and proxy model, where neural networks are used frequently. Learned proxy models or classifiers are particularly prone to incorrect predictions because they are trained on a different data distribution from the model under test (Quiñonero-Candela et al., 2008; Baek et al., 2022). Similarly, decreased model complexity may result in poor predictions (Bargagli Stoffi et al., 2022; Hu et al., 2021; Lee et al., 2020). As such, several works we have listed in Table 3 use LVLM-based evaluators (e.g., GPT-4) to increase evaluation model complexity and cover a broader data distribution. However, as we have found throughout this work, LVLMs should not be overly relied upon for consistently accurate estimates. Instead, learned hallucination detection models that further output a calibrated confidence score (relaying their uncertainty of a prediction) can assist designers with choosing how to utilize the evaluated model’s decision. Additionally, classic machine metrics like accuracy, precision, recall, false-positive rate (FPR), etc. provide engineers with an understanding of the effectiveness of hallucination detection methods prior to deployment. The learned intervention algorithm should be tuned to balance true-positive rate (TPR) and FPR, such that it does not miss critical hallucinations, nor act overly conservative (predicting hallucinations too frequently). We also suggest caution when using learning-based similarity metrics like BERTScore, SummaC, AlignScore, CLIP Score, and others listed in Table 3 for similar reasons. Specifically, while these metrics have been shown to reasonably reflect the semantic similarity of varied terms, there are still cases of poor alignment (Hanna and Bojar, 2021; Hessel et al., 2021). Even adversarial models used in adversarial prompting approaches are not safe from learned biases, which could lead to a poor understanding of the uncertainty of the model under test. For example, an adversarial agent tasked with prompting an autonomous driving LVLM could provide incorrect sensor readings, obviously resulting in poor decisions from the LVLM under test. In this case, we have not learned any additional information on the reliability of the predicted action. Thus, we argue that deep-learning-based adversarial prompters should be grounded in accurate data to truly understand the uncertainty of model predictions. Overall, engineers should take care when using deep learning methods for hallucination intervention — or as metrics during evaluation — because of their tendency to act unpredictably with out-of-domain data, and limited theoretical guarantees.